13) A nurse hangs a bag of D51/2 NS with 20meq kel. The bag is a 1 liter bag and needs to be infused at 125ml/hour. She needs to deliver a bolus of fluid first of 250 ml to be infused in 30 minutes before she can run the regular IV rate. She has an IV set that is 10gtts/ml. what rate in drops /minute will the Bolus have to run?
Answer #1

Fluid rate to administer within 1 hour =125 ml/hour

Drops per ml=10

Drip rate =(volume(ml)×drops)/(ml/time×60)

=(125×10)/(1×60)=1250/60=20.83

That is drip rate = 21 drops/minute